# Extremal metrics involving scalar curvature

### v3

@@ 5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
Are there any manifolds of non-positive Yamabe invariants that satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6?
@@ 6.1 Other rigidity results
6.2 A sufficient condition for a metric to be a positive Yamabe metric
Theorem 6.2 ([14]).
$R_{g}\cdot g\geq R_{h}\cdot h~~\mathrm{on}~M.$ (24)
On the other hand, according to [1], if $g$ is a strongly stable unique non-negative Yamabe metric with unit volume in its conformal class (then every metric sufficiently $C^{\infty}$-close to $g$ also contains the unique Yamabe metric in its conformal class), then for any $(0,2)$-tensor $h$ with $\mathrm{tr}_{g}h=0$, we have
for all sufficiently small $t>0$. Here, $\gamma_{t}\in[g_{t}:=g+th]$ is the unique Yamabe metric in its conformal class with unit volume. Since $R_{g}=\mathrm{const}$, $h=-\overset{\circ}{\mathrm{Ric}}_{g}$ is a tt-tensor with respect to $g$. Moreover, since $\gamma_{t}$ is the unique Yamabe (hence constant scalar curvature) metric in $\gamma_{t}]$ for all sufficiently small $t>0$,
$\gamma_{t}-g=t\,\mathrm{pr}_{TT}(h)+o(t)=th+o(t),$
where $\mathrm{pr}_{TT}$ is the projection onto the subspace consisting of all tt-tensors in $T_{g}\mathcal{M}_{1}$. Here, $\mathcal{M}_{1}$ is the space of all Riemannain metric on $M$ with unit volume and $T_{g}\mathcal{M}_{1}$ is the tangent space of $\mathcal{M}_{1}$ at the metric $g$. Therefore, we have
$\begin{split}R_{\gamma_{t}}\cdot\gamma_{t}-R_{g}\cdot g&=t\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}g-(R_{g}+O(t))\,\overset{\circ}{\mathrm{Ric}}_{g}\right)+o(t)\\ &=t\left(\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}+\frac{R^{2}_{g}}{n}\right)g-R_{g}\cdot\mathrm{Ric}_{g}\right)+o(t).\end{split}$
From the above, we ask the following question.
@@ 6.2 A sufficient condition for a metric to be a positive Yamabe metric
By [25, Corollary 2], for every closed manifold $M$ of dimension $\geq 3$ with positive Yamabe invariant, there is a non-Ricci-flat scalar-flat metric $g$ on $M$. Thus, the above argument with $h=\mathrm{Ric}_{g}$ for such $g$ implies that $R_{\gamma_{t}}>0$ for any sufficiently small $t>0$ where $\gamma_{t}\in[g+t\,\mathrm{Ric}_{g}]$ is the unique unit-volume Yamabe metric in its conformal class. Therefore, in particular, such a metric $g$ is not a local maximizer of the functional : $\mathcal{M}_{1}\ni h\mapsto Y(M,[h])$.
6.3 For singular metrics
Question 6.4.
@@ 6.4 About our assumptions
In our theorems 1.3, 1.4, 1.5 and 1.6, can we weaken the assumption that $\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$” to that “the metric $g$ is not an Einstein metric”?
Question 6.6.

### v4

@@ 5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
Are there any manifolds of non-positive Yamabe invariants that satisfies the assumptions of Theorems 1.5, 1.6, 1.7 and 1.8?
5.4 Unstable csc metrics
  • (On $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$) Consider $(\mathbb{S}^{1}\times\mathbb{S}^{n-1},g_{r}=r^{2}dt^{2}+g_{std})~(r>0)$. Then, $\lambda_{1}(g_{r})=\min\{r^{-2},n-1\}$ and $R_{g_{r}}=(n-1)(n-2)$. So, $g_{r}$ is stable csc metric iff $r^{-1}\geq\sqrt{n-2}$. And we have
    $(\overset{\circ}{\mathrm{Ric}}_{g})_{ij}=\begin{cases}-\frac{(n-1)(n-2)}{n}r^{2}dt^{2}&i=j=0\\ \frac{(n-2)}{n}(g_{std})_{ij}&1\leq i,j\leq n-1\\ 0&\mathrm{otherwise}.\end{cases}$
    Here, $g_{std}$ denotes the standard spherical metric on $\mathbb{S}^{n-1}$ with constant curvature $1$. Hence, $\|\overset{\circ}{\mathrm{Ric}}_{g_{r}}\|^{2}_{g_{r}}=\frac{(n-1)(n-2)^{2}}{n}$, $\max\overset{\circ}{\mathrm{Ric}}_{g_{r}}=\frac{n-2}{n}$. Therefore, $\|\overset{\circ}{\mathrm{Ric}}_{g_{r}}\|^{2}_{g_{r}}-R_{g_{r}}\cdot\max\overset{\circ}{\mathrm{Ric}}_{g_{r}}=0$ for all $r>0$.
  • (On $\mathbb{S}^{2}\times\mathbb{S}^{2}$) Consider $(\mathbb{S}^{2}\times\mathbb{S}^{2},g_{r}=r^{2}g_{std}+g_{std})$. $\lambda_{1}(g_{r})=\min\{2r^{-2},2\}$, $\frac{R_{g_{r}}}{n-1}=\frac{2r^{-2}+2}{3}$. So, $g_{r}$ is stable csc iff $1/\sqrt{2}\leq r\leq\sqrt{2}$. And we have
    $(\overset{\circ}{\mathrm{Ric}}_{g})_{ij}=\begin{cases}\left(\frac{1}{2}-\frac{r^{2}}{2}\right)(g_{std})_{ij}&i,j=1,2\\ \left(\frac{1}{2}-\frac{r^{-2}}{2}\right)(g_{std})_{ij}&i.j=3,4.\end{cases}$
    Here, $g_{std}$ denotes the standard spherical metric on $\mathbb{S}^{2}$ with constant curvature $1$. Hence, $\|\overset{\circ}{\mathrm{Ric}}_{g_{r}}\|^{2}_{g_{r}}=(1-r^{-2})^{2}$, $R_{g_{r}}=2r^{-2}+2$ and $\max\overset{\circ}{\mathrm{Ric}}_{g_{r}}=\max\left\{\frac{r^{-2}}{2}-\frac{1}{2},\frac{1}{2}-\frac{r^{-2}}{2}\right\}$. Therefore,
    $\|\overset{\circ}{\mathrm{Ric}}_{g_{r}}\|_{g_{r}}^{2}-R_{g_{r}}\cdot\max\overset{\circ}{\mathrm{Ric}}_{g_{r}}=\begin{cases}-2r^{-2}(1-r^{-2})&r\geq 1\\ 2(1-r^{-2})&r\leq 1.\end{cases}$
Unfortunately, all the metrics in these two examples do not satisfy the assumption of Theorem 1.7 with $f=0$.
Question 5.2.
For each unstable csc metric in these two examples, does a non-constant function $f$ exist that satisfies the assumption of Theorem 1.7?
@@ 6.1 Other rigidity results
6.2 For singular metrics
@@ 6.3 About our assumptions
Question 6.4.
In our theorems 1.5, 1.6, 1.7 and 1.8, can we weaken the assumption that $\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$” to that “the metric $g$ is not an Einstein metric”?
@@
According to Theorem 1.7, we propose the following problem.
Problem 6.1.
For a metric $g$ whose scalar curvature is constant, study the image of the map
$\mathcal{TT}_{g}\ni h\mapsto\langle\mathrm{Ric}_{g},h\rangle_{g}\,g-\langle\mathrm{Ric}_{g},g\rangle_{g}\,h\in\mathcal{S}^{2}TM$
where $\mathcal{TT}_{g}$ denotes the space of all $(0,2)$-tensors that are traceless and divergence-free with respect to $g$. In particular, investigate when it becomes positive or negative definite.